Sodium bromate reacts with calcium


What is the percentage yield if 82.7 g of sodium bromate reacts and 62.55 g of the calcium product is experimentally produced? 


1
Expert's answer
2021-04-13T02:07:03-0400

NaBrO3 + Cl2

Molar Mass of NaBrO3 = 150.89

= 82.7/150.89 = 0.55

= (0.55 × 62.55)/82.7= 0.4145×100

= 41.45%